OpenStudy (anonymous):

Will reward medal (if told how). Prove the identity: tan^2(x) - sin^2(x) = tan^2(x) * sin^2(x) using the following "reasons". Step 1: "This follows the definition of tanx=sinx/cosx." Step 2: "This follows from factoring out sin^2(x)." Step 3: "This follows by the definition sec x = 1/cos x." Step 4: "This follows from the Pythagorean identity sec^2(x) - 1 = tan^2(x)"

OpenStudy (anonymous):

I'm pretty sure I have step 1 down, but I can't seem to get step 2. Any help would be greatly appreciated. :)

OpenStudy (debbieg):

For step 1, you re-wrote the LHS as: \[\frac{ \sin^2(x) }{ \cos^2(x) } - \sin^2(x)\] right? Now factor out the \(\sin^2(x)\) in that expression. What do you get now on the LHS?

OpenStudy (anonymous):

Sorry, but I can't see the math that you're writing. It's saying "[Math processing error]". I rewrote the LHS as (sin^2(x)/cos^2x) - sin^2(x)

OpenStudy (debbieg):

weird.... ok, that's what I wrote. :) No, if you factor out the sin^2(x) from that, what do you get?

OpenStudy (debbieg):

E.g., your "reasons" are basically telling you exactly what to do in each step. When you are finished, you will have gone from the LHS to the RHS.

OpenStudy (anonymous):

That's what I'm a bit confused on. Would the sin values cancel out, or do you have to find the common denominator?

OpenStudy (debbieg):

No, neither. Just factor. If I have: a/b-a that is really a/b - a/1 right? so I can factor out like so: a(1/b - 1)

OpenStudy (anonymous):

So sin^2x(1/cos^2x - 1)?

OpenStudy (debbieg):

right - good! Do you see where to go next?

OpenStudy (anonymous):

Yes, I believe so. Thank you so much! How do I give you a medal? (sorry, I'm new to the website)

OpenStudy (debbieg):

I see you figured it out. :) Thank you, and you're welcome. happy to help :)